How many lbs of Brazilian coffee costing $6.00 per lb must be mixed with how many lbs of columbian coffee costing $4.00 per lb to produce 36 lbs of a mixture costing $4.50 per lb?
Answer by stanbon(75887)

How many lbs of Brazilian coffee costing $6.00 per lb must be mixed with how many lbs of columbian coffee costing $4.00 per lb to produce 36 lbs of a mixture costing $4.50 per lb?
-------------------------------
$6.00 coffee DATA:
Let amount be "x" lbs. ; Value is 6x dollars
-----------------------------
$4.00 Coffee DATA:
Amount is "36-x" lbs. ; Value is 4(35-x) = 140-4x dollars
------------------------------
Mixture DATA:
Amount is 36 lbs : Value is 4.5(36) = 162 dollars
----------------------------
EQUATION:
value + value = value
6x + 140-4x = 162
2x = 22
x = 11 lbs (Amount of $6 coffee in the mixtue)
36-11 = 25 lbs (Amount of $4.00 coffee in the mixture)
=================
Cheers,
Stan H.